I am having problems with this question. an explanation would be awesome. i think i'm just havin problems with the fraction part.

Find the equation of the line, in standard form, with a slope of 1/4 passing through the point (9,2)

Answer by THANApHD(104)

think about the standard form of the linear equations

y=mx+c where y & x are variables; m is slope and c is Y intercept
so substitude the datas in this equation

y=(1/4)x+c
2=1/4*9 + c
c = 2- (9/4)
c = (8-9)/4
c = -1/4
so we get the standard form
y=(1/4)x-(1/4)